OpenStudy (chycora):

2. (1 pt) What is the solution to the equation ? c =

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Thinking of PEMDAS, first we'd need to distribute the 3/4 into the (4c + 16). What do you think this is? \[\frac{ 3 }{ 4 }(4c + 16)\]

OpenStudy (anonymous):

First distribute 3/4.. 3/4 * 4c 3/4 * 16

OpenStudy (anonymous):

@iGreen I've got this one if you want to go help other people :)

OpenStudy (anonymous):

Yep, just noticed :P

OpenStudy (chycora):

brb

OpenStudy (chycora):

ok back

OpenStudy (anonymous):

Okay, so, if we do: \[\frac{ 3 }{ 4 }(4c + 16)\] What do we get?

OpenStudy (chycora):

3c +12

OpenStudy (chycora):

@Nibby

OpenStudy (anonymous):

Yes, that is correct. Sorry, was doing something. So, now our equation is:\[3c + 12 = 2c + 9\] I assume we're solving for C?

OpenStudy (chycora):

k

OpenStudy (chycora):

so will i do 2c +3c @Nibby

OpenStudy (anonymous):

Well, we need to get C by itself. So, we'd do:\[3c + 12 - 9 = 2c + 9 - 9\] Which would give us:\[3c + 3 = 2c\]

OpenStudy (anonymous):

Now, to get C by itself we need to subtract 3c from both sides, which would give us what?

OpenStudy (chycora):

1

OpenStudy (anonymous):

Well, no. If we subtracted 3c from both sides, we'd get: \[3c + 3 - 3c = 2c - 3c\] Which comes to: \[3 = -1c\]

OpenStudy (anonymous):

Now we'd divide -1 from both sides to get C by itself, giving us what?

OpenStudy (chycora):

-0.33

OpenStudy (anonymous):

No, no. \[\frac{ 3 }{ -1 } = \frac{ -1c }{ -1 }\] Which would give us: \[-3 = c\] So, your answer is: \[c = -3\]

OpenStudy (anonymous):

Do you understand how I got this?

OpenStudy (chycora):

yes kn i do

OpenStudy (anonymous):

You're welcome! Make sure to close the question if everything has been answered.